# Arrow-Debreu Meets Kyle: Price Discovery Across Derivatives

### v34

@@ Abstract
We analyze price discovery in a model where an agent has arbitrary private information about state probabilities and trades state-contingent claims. Our model integrates the seminal frameworks of Arrow and Debreu (1954) and Kyle (1985). In an equivalent options formulation, the informed agent has arbitrary information about an underlying asset’s payoff distribution and trades option portfolios. We characterize the informed demand and information dynamics across markets. Our results offer the first equilibrium-based explanations for longstanding empirical practices and phenomena in option markets, including common trading patterns and the volatility smile across option strikes.
@@ 1 Introduction
A fundamental economic function of prices is to convey information. As informed agents trade to exploit their information about an asset’s expected payoff, that information is incorporated into the asset price. More broadly, agents often have information about an asset’s payoff beyond its mean, such as volatility, and to leverage such information they must trade derivatives. Indeed, a principal utility of derivatives is to offer investors targeted exposure to various aspects of the underlying asset’s payoff distribution. Such derivatives trading has long been standard practice, as covered in popular textbooks such as Hull (2015). This implies that a wide range of information beyond the expected payoff is incorporated into derivatives prices.
Yet, despite their practical importance, the existing literature lacks a basic model that articulates how derivative markets systematically incorporate diverse information signals into prices. This gap stems largely from the challenges of modeling the complex information dynamics across derivatives. We address this gap by developing a model that fully captures the nonlinearity of payoff distributions and the interdependence of trades and prices across derivatives. Integrating the foundational frameworks of Arrow and Debreu (1954) and Kyle (1985), we obtain general characterizations of the informed demand, price impact, and information efficiency of prices across derivative markets.
@@
Replicating derivatives, such as options, using AD securities provides equivalent formulations of our model that account for the structural linkages across derivatives. In an equivalent options formulation, the informed trader has private information about an underlying asset’s payoff distribution and trades options, with no restriction on the payoff distribution or his information.
The informed demand formula we obtain is simple and practical. It constructs the informed trader’s contingent-claim portfolio by assigning a positive portfolio weight to the true payoff distribution and negative weights to alternative distributions. Translated to options, the formula prescribes option strategies for trading on any given aspect of the underlying payoff distribution. It encompasses observed options trading practices within a comprehensive equilibrium framework—the first such result in the literature—while also delivering practical insights for designing option strategies.
@@
The joint supply schedules for all AD securities constitute a pricing kernel that prices all derivatives simultaneously under asymmetric information. While derivatives pricing has been extensively studied from the hedging perspective since Black and Scholes (1973), we show that informed trading explains common derivatives trading patterns and price behaviors that the hedging motive cannot account for. In particular, by extending beyond the limitations of complete-information hedging models, we show that cross-strike information dynamics explain the “volatility smile,” a well-documented empirical regularity in option-implied volatilities that reduced-form models derived from Black and Scholes (1973) have yet to satisfactorily capture.111 Large classes of reduced-form models have sought to capture the volatility smile by adjusting various specifications within the Black and Scholes (1973)-type framework. Prominent approaches include the jump-diffusion models (Merton (1976), Bates (1991)), stochastic volatility models (Heston (1993), Bates (1996), Duffie et al. (2000), Britten-Jones and Neuberger (2000), Bates (2000), Aït-Sahalia and Kimmel (2007), Christoffersen et al. (2010),Aït-Sahalia et al. (2021)), and local volatility models (Dupire (1994), Berestycki et al. (2002), and Carr and Cousot (2012)).
While empirical evidence (e.g., Roll et al. (2010)) suggests derivatives improve market efficiency, theoretical explanations are limited. Our model offers a clear explanation: the linear equilibrium of Kyle (1985) is not robust with respect to the introduction of derivatives, as the absence of these markets constrains the informed trader to suboptimal linear demands. Derivatives allow the informed trader to optimally reduce his demand for low-probability states (see Example 7.1). By incorporating information that cannot be reflected in linear trades of the underlying asset, derivatives improve market efficiency.
@@ 7.1 Informed Demand
The AD informed demand (35), in its canonical form as per Corollary 6.3, shares a similar intuition with the single-asset informed demand (5):
@@
  • Information Intensity ($\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ vs. ${\bf L}$): Both $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ and ${\bf L}$ quantify the insider’s information intensity in their respective settings. In the single-asset setting, a larger $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ means the market maker can more effectively infer the variation in the asset payoff from the variation in order flow, thus reducing informed demand. For contingent claims, a larger ${\bf L}$ means the market maker can better infer the variations in the payoff distributions across signals from the variations in cross-market order flow, which reduces informed demand.
  • Signal Influence ($(s-v_{0})$ vs. $\alpha^{*}{\bf Q}$): In the single-asset setting, $(s-v_{0})$ scales an initial one-share portfolio based on the signal’s deviation from the market maker’s prior mean. For contingent claims, $\alpha^{*}$ scales the initial long-short portfolio of ${\bf Q}$ according to the market maker’s prior uncertainty regarding the signal.
@@ 9 Conclusion
In this paper, we unify the foundational elements of Arrow and Debreu (1954) and Kyle (1985) to develop a general model of joint price discovery across contingent claim markets. By connecting observed market practices with theoretical insights, our results bridge a longstanding gap between theory and empirics. Beyond its immediate empirical implications, our basic model offers a versatile platform for further extensions. For example, in a setting with risk-averse agents, where the contingent-claim prices aggregate risk as well as information, one can explore how information asymmetry distorts risk-sharing incentives across derivative markets. Extending our model to a dynamic environment provides a framework to study the temporal evolution of information spillovers across markets. These present pathways for future research.

### v35

@@ Abstract
We study price discovery in a model where an informed agent has arbitrary private information about state probabilities and trades state-contingent claims. The model unifies the seminal frameworks of Arrow and Debreu (1954) and Kyle (1985). When the claims are options, the informed agent has arbitrary information about the underlying asset’s payoff distribution and trades option portfolios. We characterize the informed demand and cross-market information dynamics. Our results provide the first equilibrium-based explanation for longstanding empirical practices and regularities in option markets, such as common trading strategies and the volatility smile across option strikes.
@@ 1 Introduction
A fundamental economic function of prices is to convey information. As informed agents trade to exploit their information about an asset’s expected payoff, that information becomes incorporated into the asset price. More broadly, agents may have information about aspects of an asset’s payoff beyond just its expected value—such as volatility—which can only be exploited through trading derivatives. This role of derivatives in providing targeted exposure to various aspects of an asset’s payoff distribution has been well understood since Ross (1976). Indeed, such derivatives trading strategies have long been standard practice, as reflected in widely used textbooks like Hull (2015). This implies that a broad range of information beyond an asset’s expected payoff is incorporated in derivatives prices.
Yet, the existing literature lacks a basic model articulating how derivative markets systematically incorporate diverse information signals into prices. This gap stems largely from the challenges of modeling the complex information dynamics across derivatives. We address this gap by developing a model that unifies the foundational frameworks of Arrow and Debreu (1954) and Kyle (1985). The model fully captures the nonlinearity of payoff distributions and the interdependence of trades and prices across derivatives. We obtain general characterizations of the informed demand, price impact, and information efficiency of prices across derivative markets.
@@
Replicating derivatives—such as options—using AD securities yields alternative formulations of our model that account for the structural linkages across derivative instruments. In an options formulation, the informed trader has arbitrary private information about an underlying asset’s payoff distribution and trades options.
The equilibrium informed demand formula is simple and practical. It constructs the informed trader’s contingent-claim portfolio by assigning a positive portfolio weight to the true payoff distribution and negative weights to alternative distributions. For options, the formula prescribes option strategies for trading on any given aspect of the underlying payoff distribution. It encompasses observed options trading practices within a comprehensive equilibrium framework—the first such result in the literature—while also delivering practical insights for designing option strategies.
@@
The joint supply schedules for all AD securities form a pricing kernel that prices all derivatives simultaneously under asymmetric information. While derivatives pricing has been extensively studied from the hedging perspective since Black and Scholes (1973), we move beyond the limitations of complete-information hedging models and show that informed trading explains observed derivative price behaviors that the hedging motive cannot account for. In particular, we show that cross-strike price formation dynamics explain the “volatility smile”—a well-documented empirical regularity in option-implied volatilities that existing reduced-form models derived from Black-Scholes do not fully capture.111 Numerous reduced-form models have attempted to capture the volatility smile by modifying components within Black-Scholes-type frameworks. Prominent approaches include the jump-diffusion models (Merton (1976), Bates (1991)), stochastic volatility models (Heston (1993), Bates (1996), Duffie et al. (2000), Britten-Jones and Neuberger (2000), Bates (2000), Aït-Sahalia and Kimmel (2007), Christoffersen et al. (2010), Aït-Sahalia et al. (2021)), and local volatility models (Dupire (1994), Berestycki et al. (2002), and Carr and Cousot (2012)).
While empirical evidence (e.g., Roll et al. (2010)) suggests derivatives improve market efficiency, theoretical explanations have been limited. Our model offers a clear explanation: the linear equilibrium of Kyle (1985) is not robust with respect to the introduction of derivatives, as the absence of these markets constrains the informed trader to suboptimal linear demands. Derivatives allow the informed trader to optimally reduce his demand for low-probability states (see Example 7.1). By incorporating information that cannot be reflected in linear trades of the underlying asset, derivatives improve market efficiency.
@@ 7.1 Informed Demand
The AD informed demand (35) extends the intuition behind the single-asset informed demand (5) to the contingent claims setting. This is made clear by comparing the AD informed demand in its canonical form (defined in Corollary 6.3) with the single-asset informed demand—both being determined by the information intensity and signal effect:
@@
  • Information Intensity ($\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ vs. ${\bf L}$): Both $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ and ${\bf L}$ characterize the insider’s information intensity in their respective settings. In the single-asset setting, a larger $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ allows the market maker to better infer the variation in the asset payoff from the fluctuations in order flow, thus reducing informed demand. For contingent claims, a larger information intensity matrix ${\bf L}$ improves the market maker’s inference about the payoff distribution from cross-market order flow, thereby dampening informed demand.
  • Signal Effect ($(s-v_{0})$ vs. $\alpha^{*}{\bf Q}$): In the single-asset setting, the signal effect is captured by $(s-v_{0})$, which scales the insider’s demand based on the signal’s deviation from the market maker’s prior estimate $v_{0}$. For contingent claims, the endogenous constant $\alpha^{*}$ scales the initial long-short portfolio of ${\bf Q}$ according to the market maker’s prior uncertainty regarding the signal.
@@ 9 Conclusion
In this paper, we unify the elements of Arrow and Debreu (1954) and Kyle (1985) to develop a general model of price discovery across contingent claim markets. By encompassing observed market practices in an equilibrium framework, our results bridge a longstanding gap between theory and empirics. Beyond its immediate empirical relevance, the model’s simplicity and generality make it adaptable for various extensions. For instance, in settings with risk-averse agents, the model allows exploration of how information asymmetry distorts risk-sharing incentives across contingent-claim markets. Extending the model to a dynamic setting offers a framework for analyzing the temporal evolution of information spillovers across markets. These extensions are new avenues for future research on derivative market behavior and cross-market information transmission.